40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  SUS Lab eLog  Not logged in ELOG logo
Message ID: 662     Entry time: Tue Jul 9 01:24:56 2013
Author: Giorgos 
Type: Summary 
Category: SUS 
Subject: HE sensors test, arrangement, and offset & Strain Gauge arrangement 

Today, we first talked about the connection of the HE conditioning boards to the HE sensors and the arrangement of the wires on the connector. There are 7 HE sensors named after their position (e.g. W=West). Starting from the right, the first four pins denote the sensors that lie above the plate.The bottom row is the bottom part of the connector to the HE signal. X denotes the pins not used and the last three pin places on the left are for the sensors imbedded in the coils, which are though--for the time being--not used.




We tested the transfer functions (TF) of our 7 HE sensor conditioning boards. Six of them had identical TF, same as the ones we expected and one of them (S1) had a similar TF, but a totally different phase. We extracted them to a floppy disk and inserted them to a computer, where we created files that contain the data of the TF plots. Tomorrow, we need to plot the data in Mathematica. We also measured the offset for our Hall-effect sensors on the oscilloscope. We used Vin=0 to measure the actual offset and then adjust R2 to null it. Here are the recorded offsets:

AC1:2.37V, AC2: not working, AC3:2.34V, S1:2.5V, S2:2.5V, N:2.5V, W:2.46V.

We also looked at the connector for the Strain Gauge (S.G.) and DC motors (M). We have six connections for each. We named our S.G. boards, depending on the location of the corresponding--in our setup--strain gauge. IMoving from the right to the left, the strain gauge sensors correspond to: TS (top south), TW, TE, BN (bottom north), BS, BE. We found a problem with the BE op-amp; it must be broken. We tested the output signal of some boards and we did not find a steady DC amplified voltage we expected; we thought of introducing a low-pass filter (since DC signals have ideally a 0Hz frequency)before the signal reaches the strain gauge op-amp.


Tomorrow, we will measured the TF of the Strain Gauge boards to see what is wrong. We will also insert a low-pass filter with a cut-off frequency around 10Hz.

Attachment 2: Strain_Gauge_&_Motors_Connections.png  6 kB  | Hide | Hide all
ELOG V3.1.3-